A biochemist studying breakdown of the insecticide DDT finds that it decomposes by a first-order reaction with a half-life of 12.0 yr. How long does it take DDT in a soil sample to decompose from 592 ppbm to 10.0 ppbm (parts per billion by mass)?
Answer #1

for first order reaction

K = 0.693 / t1/2

K = 0.693 / 12.0

K = 0.058 yr-1

K = (2.303 /t) log [A0] / [A]

[A0] = 592 ppbm

[A] = 10.0 ppbm

K = 0.058 yr-1

t = ?

0.058 = (2.303 /t) log [592] / [10.0]

0.058 = (2.303 /t) 1.77

2.303 /t = 0.033

t = 69.79 yr